OpenStudy (anonymous):

What is the exact value of the expression √12 + √50 - √48? Simplify completely.

OpenStudy (anonymous):

@jim_thompson5910 Thanks again. :)

OpenStudy (anonymous):

@ParthKohli Help?

Parth (parthkohli):

Do you know radicals?

OpenStudy (anonymous):

Yes a little im just not totally sure on how to do them.

Parth (parthkohli):

You just factor. For example \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{12} = \sqrt{\cancel{2 * 2} * 3} = 2\sqrt3 }\)

OpenStudy (anonymous):

Ok what about the 50...

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow \sqrt{50} = \sqrt{5 * 2 * 5} = \sqrt{\cancel{5 * 5} * 2} = {5\sqrt{2}} }\)

Parth (parthkohli):

@lgbasallote link your tutorial please

OpenStudy (lgbasallote):

nahh it was a flop :p

OpenStudy (anonymous):

How do you put everything together?

Parth (parthkohli):

For example: \(\Large \color{MidnightBlue}{\Rightarrow 2\sqrt{2} - 3\sqrt2 = -1\sqrt2 }\) or \(\Large \color{MidnightBlue}{\Rightarrow 19\sqrt{13} + 14\sqrt{13} = 33\sqrt{13} }\) But, you can't do this: \(\Large \color{MidnightBlue}{\Rightarrow 4\sqrt{y} + 8\sqrt{x} = 12\sqrt{xy} }\)

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

I still dont know the full answer to my question. :/

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow \sqrt{48} = \sqrt{2 * 2 * 2 * 2 * 3} = 4\sqrt{3} }\)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 2\sqrt{3} + 5\sqrt{2} - 4\sqrt{3} = -2\sqrt3 + 5\sqrt2 }\)

OpenStudy (anonymous):

ok thanks